# Efficient Mixture-of-Expert for Video-based Driver State and Physiological Multi-task Estimation in Conditional Autonomous Driving

### v1

@@ III-C2 VDMoE Block
To learn the underlying correlations between different tasks, based on the classic MoE block [36], we proposed the hierarchical VDMoE architecture to accommodate the spatio-temporal characteristics of heterogenous facial features. There are $L$ VDMoE blocks in total. Specifically, as previous feature embedding modules have already encoded heterogenous facial features to high-dimension space, to minimize the computational cost and avoid the decline in generalizability due to high complexity, we instantiated each expert in the VDMoE block with a two-layered MLP. Additionally, given our $m$ put the spatial information from heterogenous facial features at the last dimension, following the matrix multiplication in linear algebra, the weights of MLP experts can only be applied to the last dimension (i.e., spatial information). However, the temporal information is also important [54], as both drivers’ behavior and physiological reflection under specific states should not be instantaneous. Thus, we introduced the spatio-temporal expert mechanism in our VDMoE block. Specifically, given that only spatial information can be processed in $m$, we first transposed the $m$ at the last two dimensions and obtained the $m_{t}\in\mathbb{R}^{V\times T}$, where temporal information of each spatial channel can be learned. Then, in each VDMoE block, we constructed $K$ experts, which can be further equally classified as spatial experts $E_{s}(*)$ and temporal experts $E_{t}(*)$ corresponding to $m$ and $m_{t}$ respectively. Experts share no parameters and are initialized separately. The structure of each type of expert is shown in Figure 3. Then, the $m$ was processed to $\frac{K}{2}\ m_{s}^{{}^{\prime}}$ by $\frac{K}{2}$ different spatial experts $E_{s}(*)$. Similarly, we obtained $\frac{K}{2}\ m_{t}^{{}^{\prime}}$ after input $m_{t}$ to $\frac{K}{2}\ E_{t}(*)$. Moreover, to avoid information degradation with the increase in the number of layers in neural networks [38], we leveraged the jump connection in the residual network. Therefore, the intermediate features $m^{{}^{\prime}}$ consist of $K/2m_{s}^{{}^{\prime}}$, $K/2m_{t}^{{}^{\prime}}$, and the initial input of the VDMoE block $m$. The whole procedure is formalized as follows:
@@
Then, we designed an expert router mechanism $R(*)$ to dynamically aggregate the above intermediate features. Each router $R(*)$ corresponds to a specific intermediate feature. The router $R(*)$ took the $m$ as the input and nonlinearly converted it from $\mathbb{R}^{T\times V}$ to a probability vector of shape $(K+1)\times 1$. In detail, upon receiving the $m$, we first applied average pooling to compress the temporal dimension. Then, a two-layer MLP performed a nonlinear transformation, followed by a Softmax function to generate the output. Finally, the probability vector generated by the router was used to selectively activate features by weighting the intermediate variables. Each element of the vector represents the activation probability for a corresponding feature. By weighted summing these intermediate features with probabilities features, we emphasized important features while suppressing less relevant ones. The aggregated feature $m^{{}^{\prime}}$ was sent to the next VDMoE block.
@@ III-C3 Task-Related Feature Gating and Estimation Head
After the VDMoE blocks, a representative task-generic feature $m^{{}^{\prime}}\in\mathbb{R}^{T\times V}$ was obtained. Although there are correlations among multiple target tasks and can be learned by one shared backbone, separate feature spaces for different tasks are still necessary [27]. Therefore, we designed element-wise gated networks to select sparse high-level representations for specific tasks. Specifically, four element-wise gates $G_{i}(*)$ were initiated, which took the original embedding feature $m$ as input. Then, seeing Figure 3, after the average pooling and a two-layered MLP, the output of each gate $G_{i}(*)$ was controlled by a Sigmoid function, which makes each element of the output vector in the range [0, 1]. We argue that, being different from Softmax-based weighted aggregation [36, 64], the element-wise Sigmoid-based gate can adaptively control the importance of low-level information relative to specific tasks, thus selecting proper subspaces $\{m^{{}^{\prime}}_{i}\}^{4}\in\mathbb{R}^{V}$ from a uniform task-generic space $m^{{}^{\prime}}$. Lastly, we inputted these subspaces into four task-specific estimation heads as:
@@ III-D Prior-driven Alignment
As has been pointed out in previous studies [25], there was an overall negative correlation between drowsiness and cognitive load. In general, the prolonged low workload environments can lead to a ‘dormant’ state, increasing drowsiness; whereas complex tasks can be cognitively demanding and thus help keep the brain active and delay the onset of drowsiness [65]. Inspired by these, we designed a prior-driven soft regularization $\mathcal{L}_{align}$ to keep a general negative relationship between the estimated distribution of drowsiness and cognitive load.
@@
III-E Optimization Goal
To facilitate the estimation of drowsiness, cognitive load, HR, and RR. We designed different optimization goals depending on the property of each task. For the loss of drowsiness $\mathcal{L}_{drow}$ and cognitive load $\mathcal{L}_{cog}$, given there is strong subjective consciousness and individual differences in assessing drowsiness and cognitive load with the self-report questionnaires [67], participants will label their current status with varied standards. When training the model with data from different individuals, the mapping between facial features and state labels is not robust. Thus, we applied the generalizable cross-entropy loss (Truncated Loss) [68] to optimize the drowsiness and cognitive load estimation, and the smooth L1 loss for the estimation of HR $\mathcal{L}_{hr}$ and RR $\mathcal{L}_{rr}$ following [60]. Besides, to suppress some meaningless effects of the regularization at the early iterations, we applied the adaptation factors $\lambda$ as Eq. (5)
$t=\frac{Iter_{current}}{Iter_{total}},$ (4)
$\lambda=\frac{2}{1+\exp{-10t}}.$
@@ IV-A3 Apparatus
Besides, based on the KSS questionnaire collected after each trial, there are 75.86% samples labeled as ’awake’ (KSS <5), and 24.14% labeled as ’drowsy’ (KSS $\geq$ 5). Similarly, using the NASA-TLX questionnaire, there are 29.61% labeled as ’high cognitive load’ (weighted score >10), and 60.39% samples labeled as ’normal cognitive load’ (weighted score $\leq$ 10).
@@ IV-B Other Datasets
We also evaluated our method in other two public datasets. However, as there is no dataset that satisfies the requirements of the multi-task remote detection (i.e., more than one driver state, and contains physiological signals), we selected one dataset for detecting drowsiness (FatigueView) [32] and one another dataset for multi-task vital signs (V4V) [75].
@@ V-A Implementation Details
We separated all subjects in each dataset into two parts: 80% for training and 20% for testing. Before training, to ensure a more accurate RR and HR measurement, we set the sliding window size $F=300$ frames and step size $S=30$ frames. Then, the MediaPipe Face Mesh package from Google [^1] https://github.com/google-ai-edge/mediapipe was leveraged to perform landmark detection. Following [17], we selected 106 facial landmarks. The size of landmark matrix $I_{f}$ is $\mathbb{R}^{300\times 106\times 2}$. To standardize the input, each frame of the left eye and right eye $I_{l},I_{r}$ were resized to $\mathbb{R}^{25\times 25\times 3}$ and the mouth frame was resized to $\mathbb{R}^{35\times 15\times 3}$. Additionally, based on 106 landmarks of each sliding window, we further generated STMaps following [78]. The size of STMap is $\mathbb{R}^{300\times 25\times 3}$. Then, in Table III, we present the specific configuration of our feature embedding modules. In practice, we set up $K=4$ experts in each VDMoE block, and there are $L=1$ layers of VDMoE. Adam optimizer with a learning rate of 0.00001 was used for training. The trade-off parameter $k$ was 0.001. The batch size $B$ and maximum iterations were set to 250 and 20000 respectively.
Following existing works [79, 20], we used Accuracy, F1 score, Sensitivity, and Specificity to evaluate performance on cognitive load and drowsiness estimation. For HR and RR estimation, mean absolute error (MAE), root mean square error (RMSE), and Pearson’s correlation coefficient (p) were used [80]. We conducted five evaluations using five different random seeds, and then used a paired t-test to determine the significance of the mean performance difference between the best and second-best models. The whole work was implemented in the Pytorch framework and all experiments were conducted on a Nvidia RTX A6000.
@@ V-B Baseline Models
To evaluate the performance of our proposed VDMoE, we compared it with several single-task or multi-task models for both state and physiological estimation. Specifically, for single-task models, DBN-HMM [29], 3DCNN+BiLSTM [56], DDDNet [57], IsoSSL-MoCo [19], ReNeXt3D-101+LSTM [58], FacialUnits [59], 2s-STGCN [54], and VBFLFF [17] were used for drowsiness estimation, and CHROM [47], POS [48], ARM-RR [49], Dual-GAN [50], and PhysFormer++ [51] were used for physiological measurement. For multi-task models, since there is currently no facial video-based method proposed for multi-task state and physiological estimation, except for three methods for multi-task physiological estimation (i.e., MTTS-CAN [52], BigSmall [53], and PhysMLE [27]), we built four baselines with the hard-parameter sharing paradigm. In general, each multi-task baseline consists of a typical backbone network (i.e., ResNet3D and ViViT for 3D video input, ResNet and ViT for physiological estimation with STMap input [60]) and estimation heads that are same as VDMoE. The results are presented in Table V, VI.
V-C Rsults of Comparison Experiment
@@ V-E Impact of $K$ and $L$
The $K$ denotes the number of experts in each layer of the model, while the $L$ denotes the number of layers in the VDMoE. Since the $K$ and $L$ control the width and depth of the VDMoE, we conducted a hyper-parameter test to verify the best choice of $K$ and $L$, and their impact on the multi-task performance. We visualized the results in the Figure 6.
@@ V-F Performance on Other Datasets
To verify the effectiveness of VDMoE in wider scenarios, we assessed it on two other public datasets (i.e., FatigueView and V4V). The results are presented in Table IX. For the multi-task remote physiological measurement, the VDMoE still significantly outperformed all methods. For the single-task drowsiness estimation in the FatigueView dataset, despite the VDMoE still outperforming most of the baselines, the ResNet3D stands out with the highest accuracy (89.37%), F1-score (93.78%), and specificity (91.30%). It suggests that the ResNet3D is particularly effective at identifying true negatives, i.e., correctly identifying awake states. However, considering the significantly larger computational cost of ResNet3D, we believe that the VDMoE shows competitive single-task performance.
@@ VI Discussion
Firstly, the VDMoE model’s outstanding performance, particularly in comparison with single-task models and multi-task baselines, underscores the effectiveness of its architecture. Notably, the model’s superior accuracy in drowsiness estimation, even when compared to specialized single-task models, suggests that a well-designed multi-task model does not necessarily trade off performance in individual tasks for versatility. Besides, the better performance of VDMoE compared to multi-task baselines without task-specific feature subspace construction (e.g., ResNet3D and ViViT) illustrated that shared low-level information extraction and separate task-related feature space can enhance performance across related tasks by leveraging common features. Moreover, the comparative underperformance of models relying solely on facial video inputs versus those utilizing preprocessed facial features highlights the potential limitations of raw video data in capturing subtle physiological or behavioral cues. This finding suggests that preprocessing steps or feature engineering might still play a crucial role in optimizing model performance, even in an era where end-to-end deep learning models are prevalent. Particularly, considering both facial features at the action level (e.g. blinks and expressions) and changes in color (STMap) during pre-processing may help extract changes in physiological indicators associated with the driver states.
In addition, the inclusion of a prior-driven loss, denoted as $\mathcal{L}_{align}$ plays a pivotal role in aligning the model’s outputs with prior knowledge. This loss function helps guide the learning process, which ensures that the model not only optimizes estimation accuracy but also considers the underlying structure or relationships within the data that are known as priori. By incorporating $\mathcal{L}_{align}$, the model is encouraged to learn representations that are consistent with the findings in previous research (i.e., the correlation between drowsiness and cognitive load), potentially leading to more robust and generalizable performance. In the future, exploring more sophisticated forms of prior-driven losses that can be explicitly guided by subtle relationships among tasks could further improve the model’s robustness and generalizability.
@@ VII Limitations
While the VDMoE model demonstrates impressive performance in multi-task driver state and physiological measurement estimation in conditional autonomous driving, some limitations still exist in the current study. Firstly, although the VDMoE achieved a balance between performance and computational efficiency through its architecture, the challenges for model interpretability are still unsolved. As a neural network with complex high-dimension mapping and nonlinear transform, understanding why the model makes specific decisions or how it differentiates between tasks at a granular level is important in building trustworthy AI systems [81].
Besides, while this study proposed new multi-task datasets in the context of driving, the generalizability of the model to real-world scenarios across different individuals, driving conditions, and sensor setups remains a concern. The potential for biases between the training and testing data could impact the model’s effectiveness and fairness when deployed in diverse real-world environments. Particularly, real-world driving conditions are dynamic, with rapid changes in external factors (e.g., light, weather, traffic conditions), which might be challenging for assessing the driver’s physiological indicators or state with the camera. The model’s ability to adapt to these changes in real-time, especially without retraining or manual adjustments, is an area that requires further exploration.

### v2

@@ III-C2 VDMoE Block
To learn the underlying correlations between different tasks, based on the classic MoE block [37], we proposed the hierarchical VDMoE architecture to accommodate the spatio-temporal characteristics of heterogenous facial features. There are $L$ VDMoE blocks in total. Specifically, as previous feature embedding modules have already encoded heterogeneous facial features to high-dimensional space, to minimize the computational cost and avoid the decline in generalizability due to high complexity, we instantiated each expert in the VDMoE block with a two-layered MLP. Additionally, given our $m$ put the spatial information from heterogenous facial features at the last dimension, following the matrix multiplication in linear algebra, the weights of MLP experts can only be applied to the last dimension (i.e., spatial information). However, the temporal information is also important [55], as both drivers’ behavior and physiological reflection under specific states should not be instantaneous. Thus, we introduced the spatio-temporal expert mechanism in our VDMoE block. Specifically, given that only spatial information can be processed in $m$, we first transposed the $m$ at the last two dimensions and obtained the $m_{t}\in\mathbb{R}^{V\times T}$, where temporal information of each spatial channel can be learned. Then, in each VDMoE block, we constructed $K$ experts, which can be further equally classified as spatial experts $E_{s}(*)$ and temporal experts $E_{t}(*)$ corresponding to $m$ and $m_{t}$ respectively. Experts share no parameters and are initialized separately. The structure of each type of expert is shown in Figure 1. Then, the $m$ was processed to $\frac{K}{2}\ m_{s}^{{}^{\prime}}$ by $\frac{K}{2}$ different spatial experts $E_{s}(*)$. Similarly, we obtained $\frac{K}{2}\ m_{t}^{{}^{\prime}}$ after input $m_{t}$ to $\frac{K}{2}E_{t}(*)$. Moreover, to avoid information degradation with the increase in the number of layers in neural networks [39], we leveraged the jump connection in the residual network. Therefore, the intermediate features $m^{{}^{\prime}}$ consist of $\frac{K}{2}\ m_{s}^{{}^{\prime}}$, $\frac{K}{2}\ m_{t}^{{}^{\prime}}$, and the initial input of the VDMoE block $m$. The whole procedure is formalized as follows:
@@
Then, we designed an expert router mechanism $R(*)$ to dynamically aggregate the above intermediate features. Each router $R(*)$ corresponds to a specific intermediate feature. The router $R(*)$ took the $m$ as the input and nonlinearly converted it from $\mathbb{R}^{T\times V}$ to a probability vector of shape $(K+1)\times 1$. In detail, upon receiving the $m$, we first applied average pooling to compress the temporal dimension. Then, a two-layer MLP performed a nonlinear transformation, followed by a Softmax function to generate the output. Finally, the probability vector generated by the router was used to selectively activate features by weighting the intermediate variables. Each element of the vector represents the activation probability for a corresponding feature. By weighted summing these intermediate features with probability features, we emphasized important features while suppressing less relevant ones. The aggregated feature $m^{{}^{\prime}}$ was sent to the next VDMoE block.
@@ III-C3 Task-Related Feature Gating and Estimation Head
After the VDMoE blocks, a representative task-generic feature $m^{{}^{\prime}}\in\mathbb{R}^{T\times V}$ was obtained. Although there are correlations among multiple target tasks and can be learned by one shared backbone, separate feature spaces for different tasks are still necessary [28]. Therefore, we designed element-wise gated networks to select sparse high-level representations for specific tasks. Specifically, four element-wise gates $G_{i}(*)$ were initiated, which took the original embedding feature $m$ as input. Then, seeing Figure 1, after the average pooling and a two-layered MLP, the output of each gate $G_{i}(*)$ was controlled by a Sigmoid function, which makes each element of the output vector in the range [0, 1]. We argue that, being different from Softmax-based weighted aggregation [37, 65], the element-wise Sigmoid-based gate can adaptively control the importance of low-level information relative to specific tasks, thus selecting proper subspaces $\{m^{{}^{\prime}}_{i}\}^{4}\in\mathbb{R}^{V}$ from a uniform task-generic space $m^{{}^{\prime}}$. Lastly, we input these subspaces into four task-specific estimation heads as:
@@ III-D Prior-driven Alignment
As noted by [26], drowsiness and cognitive load tend to exhibit an overall negative correlation in driving contexts. This inverse relationship is rooted in cognitive-resources theory [66], which indicates the limited human attentional capacity: when task demands are low, surplus resources ‘wander’, leading to mind-wandering and increased drowsiness [67]. Conversely, moderate task demands optimally engage attentional control mechanisms, sustaining cortical arousal, maintaining vigilance, and delaying the onset of fatigue [68]. Studies of vigilance decrement further demonstrate that under-stimulation produces lapses in attention and microsleeps [69], whereas appropriately challenging tasks improve sustained performance. We therefore leverage this well-established theoretical and empirical foundation to introduce a prior-driven soft regularization $\mathcal{L}_{align}$, enforcing a general negative alignment between our drowsiness and cognitive workload estimates
@@
III-E Optimization Goal
To facilitate the estimation of drowsiness, cognitive load, HR, and RR. We designed different optimization goals depending on the properties of each task. For the loss of drowsiness $\mathcal{L}_{drow}$ and cognitive load $\mathcal{L}_{cog}$, given there is strong subjective consciousness and individual differences in assessing drowsiness and cognitive load with the self-report questionnaires [71], participants will label their current status with varied standards. When training the model with data from different individuals, the mapping between facial features and state labels is not robust. Thus, we applied the generalizable cross-entropy loss (Truncated Loss) [72] to optimize the drowsiness and cognitive load estimation, and the smooth L1 loss for the estimation of HR $\mathcal{L}_{hr}$ and RR $\mathcal{L}_{rr}$ following [61]. Besides, to suppress some meaningless effects of the regularization at the early iterations, we applied the adaptation factors $\lambda$ as Eq. (4)
$t=\frac{Iter_{current}}{Iter_{total}},$ (4)
$\lambda=\frac{2}{1+\exp{-10t}}.$
@@ IV-A3 Apparatus
Besides, following [79], based on the KSS questionnaire collected after each trial, there are 75.86% samples labeled as ’awake’ (KSS <5), and 24.14% labeled as ‘drowsy’ (KSS $\geq$ 8). For cognitive workload, we first normalized each participant’s raw NASA-TLX scores to the 1–20 range to mitigate inter-subject differences [80]. We then dichotomized the normalized scores at a global midpoint (10), yielding 29.61% of trials labeled as ‘high cognitive load’ (score >10) and 60.39% as ‘normal cognitive load’ (score $\leq$10). A sensitivity analysis shifting the threshold by ±2 points showed model performance variance below 5%, confirming that our median-split strategy produces robust class balances without unduly biasing results.
@@ IV-B Other Datasets
We also evaluated our method in two other public datasets. However, as there is no dataset that satisfies the requirements of the multi-task remote detection (i.e., more than one driver state, and contains physiological signals), we selected one dataset for detecting drowsiness (FatigueView) [33] and one another dataset for multi-task vital signs (V4V) [81].
@@ V-A Implementation Details
For the experiments on three datasets, the cross-subject evaluation protocol was utilized. Specifically, for all subjects in each dataset, we randomly separated them into training, validation, and test sets with the ratio 6:2:2. Models were trained on the data from subjects in the training set, adjusting parameters in the validation set, and only reporting the results in the test set. Before training, to ensure a more accurate RR and HR measurement, we set the sliding window size $F=300$ frames and step size $S=30$ frames. Then, the MediaPipe Face Mesh package from Google [^1] https://github.com/google-ai-edge/mediapipe was leveraged to perform landmark detection. Following [18], we selected 106 facial landmarks. The size of landmark matrix $I_{f}$ is $\mathbb{R}^{300\times 106\times 2}$. To standardize the input, each frame of the left eye and right eye $I_{l},I_{r}$ was resized to $\mathbb{R}^{25\times 25\times 3}$ and the mouth frame was resized to $\mathbb{R}^{35\times 15\times 3}$. Additionally, based on 106 landmarks of each sliding window, we further generated STMaps following [85]. The size of STMap is $\mathbb{R}^{300\times 25\times 3}$. Then, in Table III, we present the specific configuration of our feature embedding modules. In practice, we set up $K=4$ experts in each VDMoE block, and there are $L=1$ layers of VDMoE. Adam optimizer with a learning rate of 0.00001 was used for training. The trade-off parameter $k$ was 0.001. The batch size $B$ and maximum iterations were set to 250 and 20000, respectively.
Following existing works [86, 21], we used Accuracy, F1 score, Sensitivity, and Specificity to evaluate performance on cognitive load and drowsiness estimation. For HR and RR estimation, mean absolute error (MAE), root mean square error (RMSE), and Pearson’s correlation coefficient (p) were used [87]. We conducted five evaluations using five different random seeds, and then used a paired t-test to determine the significance of the mean performance difference between the best and second-best models. The whole work was implemented in the Pytorch framework and all experiments were conducted on an Nvidia RTX A6000.
@@ V-B Baseline Models
To evaluate the performance of our proposed VDMoE, we compared it with several single-task or multi-task models for both state and physiological estimation. Specifically, for single-task models, DBN-HMM [30], 3DCNN+BiLSTM [57], DDDNet [58], IsoSSL-MoCo [20], ReNeXt3D-101+LSTM [59], FacialUnits [60], 2s-STGCN [55], MIGCN [82], and VBFLFF [18] were used for drowsiness estimation, and CHROM [48], POS [49], ARM-RR [50], Dual-GAN [51], ConDiff-rPPG [62], and PhysFormer++ [52] were used for physiological measurement. For multi-task models, since there is currently no facial video-based method proposed for multi-task state and physiological estimation, except for three methods for multi-task physiological estimation (i.e., MTTS-CAN [53], BigSmall [54], and PhysMLE [28]), we built four baselines with the hard-parameter sharing paradigm. In general, each multi-task baseline consists of a typical backbone network (i.e., ResNet3D and ViViT for 3D video input, ResNet and ViT for physiological estimation with STMap input [61]) and estimation heads that are the same as VDMoE.
Among the baselines, following their source papers, DBN-HMM, 3DCNN+BiLSTM, DDDNet, IsoSSL-MoCo, ReNeXt3D-101+LSTM, MIGCN, CHROM, POS, ARM-RR, PhysFormer++, MTTS-CAN, BigSmall, ResNet3D, and ViViT utilize only raw 3D facial video as input. In contrast, FacialUnits and 2s-STGCN rely on facial landmarks extracted from the facial video. Dual-GAN, ConDiff-rPPG, and PhysMLE use STMap as input, while VBFLFF incorporates landmarks along with videos of the eye and mouth subregions. The results are presented in Table V, VI.
V-C Results of Comparison Experiment
@@ V-E Impact of $K$ and $L$
The $K$ denotes the number of experts in each layer of the model, while the $L$ denotes the number of layers in the VDMoE. Since the $K$ and $L$ control the width and depth of the VDMoE, we conducted a hyper-parameter test to verify the best choice of $K$ and $L$, and their impact on the multi-task performance. We visualized the results in Figure 6.
@@ V-F Performance on Other Datasets
To verify the effectiveness of VDMoE in wider scenarios, we assessed it on two other public datasets (i.e., FatigueView and V4V). The results are presented in Table VIII, IX. For the multi-task remote physiological measurement, the VDMoE still significantly outperformed all methods. For the single-task drowsiness estimation in the FatigueView dataset, despite the VDMoE still outperforming most of the baselines, the ResNet3D stands out with the highest accuracy (89.37%), F1-score (93.78%), and specificity (91.30%). It suggests that the ResNet3D is particularly effective at identifying true negatives, i.e., correctly identifying awake states. However, considering the significantly larger computational cost of ResNet3D, we believe that the VDMoE shows competitive single-task performance.
@@ VI Discussion
Firstly, the VDMoE model’s outstanding performance, particularly in comparison with single-task models and multi-task baselines, underscores the effectiveness of its architecture. Notably, the model’s superior accuracy in drowsiness estimation, even when compared to specialized single-task models, suggests that a well-designed multi-task model does not necessarily trade off performance in individual tasks for versatility. Besides, the better performance of VDMoE compared to multi-task baselines without task-specific feature subspace construction (e.g., ResNet3D and ViViT) illustrates that shared low-level information extraction and separate task-related feature space can enhance performance across related tasks by leveraging common features. Moreover, the comparative underperformance of models relying solely on facial video inputs versus those utilizing preprocessed facial features highlights the potential limitations of raw video data in capturing subtle physiological or behavioral cues. This finding suggests that preprocessing steps or feature engineering might still play a crucial role in optimizing model performance, even in an era where end-to-end deep learning models are prevalent. Particularly, considering both facial features at the action level (e.g. blinks and expressions) and changes in color (STMap) during pre-processing may help extract changes in physiological indicators associated with the driver states.
In addition, the inclusion of a prior-driven loss, denoted as $\mathcal{L}_{align}$ plays a pivotal role in aligning the model’s outputs with prior knowledge. This loss function helps guide the learning process, which ensures that the model not only optimizes estimation accuracy but also considers the underlying structure or relationships within the data that are known a priori. By incorporating $\mathcal{L}_{align}$, the model is encouraged to learn representations that are consistent with the findings in previous research (i.e., the correlation between drowsiness and cognitive load), potentially leading to more robust and generalizable performance. In the future, exploring more sophisticated forms of prior-driven losses that can be explicitly guided by subtle relationships among tasks could further improve the model’s robustness and generalizability.
@@ VII Limitations
While the VDMoE model demonstrates impressive performance in multi-task driver state and physiological measurement estimation in conditional autonomous driving, some limitations still exist in the current study. Firstly, although the VDMoE achieved a balance between performance and computational efficiency through its architecture, the challenges for model interpretability are still unsolved. As a neural network with complex high-dimensional mapping and nonlinear transform, understanding why the model makes specific decisions or how it differentiates between tasks at a granular level is important in building trustworthy AI systems [88].
Secondly, while this study proposed new multi-task datasets in the context of driving, the generalizability of the model to real-world scenarios across different individuals, driving conditions, and sensor setups remains a concern. The potential for biases between the training and testing data could impact the model’s effectiveness and fairness when deployed in diverse real-world environments. Particularly, real-world driving conditions are dynamic, with rapid changes in external factors (e.g., light, weather, traffic conditions, occlusions), which might be challenging for assessing the driver’s physiological indicators or state with the camera. The model’s ability to adapt to these changes in real-time, especially without retraining or manual adjustments, is an area that requires further exploration. Besides, in the MCDD dataset proposed in this paper, limited by the sample size, we are unable to consider the above factors. Meanwhile, there is currently a lack of a multi-task dataset that takes into account multiple challenges in a real environment simultaneously. Therefore, we expect the collection of a more complete dataset and the validation of the models on the dataset as a future work
Lastly, in this paper, the way we label the cognitive load level is a discretization method designed by ourselves based on the NASA-TLX questionnaire. Although this approach demonstrates a certain degree of robustness in terms of performance, it still lacks theoretical support. Therefore, we also expect future work exploring more solid discretization criteria for cognitive load categorization based on subjective questionnaires.